# Gravitational wave probes of particle dark matter: a review

### v4

@@ IX.1 Dephasing due to dynamical friction of intermediate mass ratio inspirals
Ref. Eda et al. (2013) first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney et al. (2022); Mitra et al. (2025). Such effects will primarily be relevant in space-based GW detectors Duque et al. (2024); Rahman et al. (2024); Chowdhuri et al. (2024); Ghodla (2025), since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller (2009).
@@
This conclusion was significantly expanded in Eda et al. (2015), in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA. This accuracy is quantified in terms of the power-law index and the masses of the two objects that are inspiraling. If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately. This is because higher values of induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
@@
Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al. (2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al. (2025). As an illustration, we show in Fig. 26 the impact of an environmental effect, the DM spike, on the SNR obtained when match-filtering with a vacuum template to analyze data in which a binary black hole system has a DM spike around it. Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo and Silk (1999). Ref. Cole et al. (2023b) argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation. However, it is important to note that in LISA, we will be in a regime in which signals of this kind have enormous SNR, and that parameter estimation and distinguishing among different GW signals will be a key challenge. Thus, it is more likely that parameter estimates will be biased rather than signals being completely missed. Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi et al. (2025). In principle, CW methods could provide quick point estimates for key binary black hole parameters, which would enable the development of a reduced set of templates that incorporate DM spikes. These templates could then be used for follow-up matched filtering searches.
@@
Figure 26: Taken from Cole et al. (2023b). Percentage SNR loss from matched filtering with a vacuum template when the injected GW signal originates from a binary black hole system surrounded by DM spikes (the injected waveform includes the effect of dynamical friction). The $x$ and $y$ axes correspond to the parameters of the DM spike profile defined in Eq. 111. The assumed observation time is one year, with component masses $m_{1}=10^{5}M_{\odot}$ and $m_{2}=10M_{\odot}$.
@@
While Ref. Cole et al. (2023b) demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref. Zwick et al. (2024) presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources. Moreover, the authors argue about the practicality of measuring such dephasing effects, noting that the early inspiral is where the effects would be most pronounced, but this is also where the GW emission is weakest; that a constant phase offset could mimic dephasing effects ; and that matched filtering over such long timescales would have practical limitations. For thee reasons.
@@
However, Ref. Wilcox et al. (2024) provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick et al. (2024). Additionally, they propose more model-agnostic approaches, such as periodic forces, that could leave characteristic signatures on the GW signal and be applied to study the evolution of the binary system.
@@
These differing viewpoints are crucial when considering the overall approach to detecting and interpreting GWs from systems with DM spikes. While Zwick et al. (2024) concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox et al. (2024) offer a promising alternative for tackling the problem. In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi et al. (2025). As mentioned in Section VII.2, CW methods could provide ways of addressing the three concerns laid out in Ref. Zwick et al. (2024) by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search. This approach can also be applied in lieu of traditional matched filtering for PBHs in ET and Cosmic Explorer.
@@ IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes
PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, similarly to IMRI systems discussed in the previous subsection, these exotic mergers could be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. As mentioned in Sections IX.1 and VII.2, one possible approach is to adapt CW methods to search for DM spikes around binary black holes. This could provide a quick, though less sensitive, way to obtain point estimates for important binary black hole parameters, such as the chirp mass. These estimates could then be used to reduce the number of templates needed to account for DM spikes, enabling a follow-up matched filtering search.
@@ IX.3 Upper limits on black-hole environments
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al. (2020).
@@ X GW probes of atomic DM
It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model. In fact, atomic DM models exist Feng et al. (2009); Kaplan et al. (2010) in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3. with a particular coupling. In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley and DiFranzo (2018); Rosenberg and Fan (2017). If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera et al. (2018). Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh et al. (2021); Abbott et al. (2022f, 2023).
@@ X.1 Constraints from observations of binary black holes
Ref. Shandera et al. (2018) first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose. If the BH merger of GW190425 Abbott et al. (2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al. (2023). In order to obtain these constraints, we note that LIGO, Virgo and KAGRA observations of binary black holes provide estimates of the chirp mass and the component masses of the system. Ref. Singh et al. (2021, 2023) jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution. These posteriors are conditioned on the output of a GW search, which is essentially predicted merger rate density that depends both on what binary black holes have been observed and power spectral density of the interferometers.
Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al. (2018); Singh et al. (2021):
@@
In Singh et al. (2021), $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott et al. (2020b) as dark binary black hole mergers.
Beyond a range of implied $M_{\rm min}$ and $m_{\chi}$, constraints can be placed on the fraction of DM taht PBHs compose as a function of the average mass of the binary black hole system. These constraints are shown in Fig. 27, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 are dark. Here, we limit ourselves to case (1) and case (3). These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$. The parameters b=[1,2],r=[2,100]b=[-1,2],r=[2,100] determine the mass distribution of dark BHs. Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions. More observations of dark BHs would be needed to further narrow this region. However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
Figure 27: Taken from Singh et al. (2021). Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA. These plots assume an atomic model of DM that forms dark BHs. The constraints are obtained in three cases: (1) assuming none of the detected BHs are dark; (2) assuming only GW190425 is a dark BH; and (3) assuming both GW190425 and GW190814 are dark BHs. The latter two cases essentially arise from a joint posterior distribution conditioned on the rate densities inferred from the search results, which also constrain the minimum allowed mass in the assumed population distribution (see text for details). Even in the absence of any dark BHs, the constraints obtained in Singh et al. (2021) significantly outperform existing microlensing experiments Allsman et al. (2001); Tisserand et al. (2007); Niikura et al. (2019) and those from dynamics of dwarf galaxies Brandt (2016); Koushiappas and Loeb (2017). Additional constraints arise from PBH searches Abbott et al. (2019); Gaggero et al. (2017); Raidal et al. (2017, 2019); Ali-Haimoud et al. (2017). SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.

### v5

@@ IX.1 Dephasing due to dynamical friction of intermediate mass ratio inspirals
Ref. Eda:2013gg first showed that the phase evolution of an intermediate-mass ratio inspiral (a binary composed to a $\mathcal{O}(1000)M_{\odot}$ and an $\mathcal{O}(1)M_{\odot}$ object) would be significantly modified at milli-hertz frequencies due to the presence of a DM spike around the intermediate-mass BH. This occurs because the less massive object sweeps through the DM spike as it orbits around the more massive one, losing energy via dynamical friction (in addition to via GWs) and speeding up the orbit Speeney:2022ryg; Mitra:2025tag. Such effects will primarily be relevant in space-based GW detectors Duque:2023seg; Rahman:2023sof; AbhishekChowdhuri:2023rfv; Ghodla:2024fit, since IMBHs inspiral and merge at much lower frequencies than stellar-mass BHs Miller:2008fi.
@@
This conclusion was significantly expanded in Eda:2014kra, in which the authors demonstrated how accurately the power-law index can be measured from a detection of an IMRI signal in LISA. This accuracy is quantified in terms of the power-law index and the masses of the two objects that are inspiraling. If the IMRI system consists of smaller primary and secondary masses, and a larger power-law index, the power-law index can be measured more accurately. This is because higher values of induce a larger phase difference, and smaller masses in the IMBH system ensures that it spends more cycles in the LISA band before merging.
@@
Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole:2022yzw, as well as distinguishing between environmental effects and modified theories of gravity Yuan:2024duo. As an illustration, we show in Fig. 26 the impact of an environmental effect, the DM spike, on the SNR obtained when match-filtering with a vacuum template to analyze data in which a binary black hole system has a DM spike around it. Here, $m_{1}=10^{5}M_{\odot}$, which was chosen to ensure that a plausible formation mechanism exists to form this kind of BH Gondolo:1999ef. Ref. Cole:2022yzw argues that a percentage SNR loss of more than 30% would compromise the prospects of detecting a (weak) signal, and shows that even smaller SNR losses would bias parameter estimation. However, it is important to note that in LISA, we will be in a regime in which signals of this kind have enormous SNR, and that parameter estimation and distinguishing among different GW signals will be a key challenge. Thus, it is more likely that parameter estimates will be biased rather than signals being completely missed. Moreover, the dephasing caused by DM spikes underscores the need to adapt CW methods, as discussed in Section III.1 and Section VI.2, to effectively search for these systems Sethi:2025ixx. In principle, CW methods could provide quick point estimates for key binary black hole parameters, which would enable the development of a reduced set of templates that incorporate DM spikes. These templates could then be used for follow-up matched filtering searches.
@@
Figure 26: Taken from Cole:2022yzw. Percentage SNR loss from matched filtering with a vacuum template when the injected GW signal originates from a binary black hole system surrounded by DM spikes (the injected waveform includes the effect of dynamical friction). The $x$ and $y$ axes correspond to the parameters of the DM spike profile defined in Eq. 111. The assumed observation time is one year, with component masses $m_{1}=10^{5}M_{\odot}$ and $m_{2}=10M_{\odot}$.
@@
While Ref. Cole:2022yzw demonstrated that distinguishing between different BH environments, including the presence of DM spikes, is feasible, Ref. Zwick:2024yzh presents a different perspective, arguing that the dephasing effects caused by these environmental factors could result from a variety of unmodeled sources. Moreover, the authors argue about the practicality of measuring such dephasing effects, noting that the early inspiral is where the effects would be most pronounced, but this is also where the GW emission is weakest; that a constant phase offset could mimic dephasing effects ; and that matched filtering over such long timescales would have practical limitations. For thee reasons.
@@
However, Ref. Wilcox:2024sqs provides a counterpoint, suggesting that such dephasing could still be tracked within a parameterized post-Einsteinian framework, offering a potential solution to the challenges raised in Zwick:2024yzh. Additionally, they propose more model-agnostic approaches, such as periodic forces, that could leave characteristic signatures on the GW signal and be applied to study the evolution of the binary system.
@@
These differing viewpoints are crucial when considering the overall approach to detecting and interpreting GWs from systems with DM spikes. While Zwick:2024yzh concerns highlight the potential difficulties in applying matched filtering to these cases, the model-agnostic approach in Wilcox:2024sqs offer a promising alternative for tackling the problem. In light of these challenges, it is clear that both traditional and alternative methods, such as the CW approaches discussed in Section III.1 and Section VI.2, should be explored in tandem to fully address the complexities involved in detecting signals from binary black holes influenced by DM spikes Sethi:2025ixx. As mentioned in Section VII.2, CW methods could provide ways of addressing the three concerns laid out in Ref. Zwick:2024yzh by providing point estimates of binary parameters that can be followed up by a sensitive and restricted matched filtering search. This approach can also be applied in lieu of traditional matched filtering for PBHs in ET and Cosmic Explorer.
@@ IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes
PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green:2020jor; Carr:2023tpt. Thus, Ref. Cole:2022ucw showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, similarly to IMRI systems discussed in the previous subsection, these exotic mergers could be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. As mentioned in Sections IX.1 and VII.2, one possible approach is to adapt CW methods to search for DM spikes around binary black holes. This could provide a quick, though less sensitive, way to obtain point estimates for important binary black hole parameters, such as the chirp mass. These estimates could then be used to reduce the number of templates needed to account for DM spikes, enabling a follow-up matched filtering search.
@@ IX.3 Upper limits on black-hole environments
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. CanevaSantoro:2023aol modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb:2016fzn as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo:1999ef; Sadeghian:2013laa. The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso:2019rou; Kavanagh:2020cfn.
@@ X GW probes of atomic DM
It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model. In fact, atomic DM models exist Feng:2009mn; Kaplan:2009de in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3. with a particular coupling. In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley:2017ttd; Rosenberg:2017qia. If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera:2018xkn. Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh:2020wiq; LIGOScientific:2021job; LIGOScientific:2022hai.
@@ X.1 Constraints from observations of binary black holes
Ref. Shandera:2018xkn first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose. If the BH merger of GW190425 Abbott:2020uma is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar:1931ih can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh:2022wvw. In order to obtain these constraints, we note that LIGO, Virgo and KAGRA observations of binary black holes provide estimates of the chirp mass and the component masses of the system. Ref. Singh:2020wiq; Singh:2022wvw jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution. These posteriors are conditioned on the output of a GW search, which is essentially predicted merger rate density that depends both on what binary black holes have been observed and power spectral density of the interferometers.
Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar:1931ih; Shandera:2018xkn; Singh:2020wiq:
@@
In Singh:2020wiq, $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott:2020khf as dark binary black hole mergers.
Beyond a range of implied $M_{\rm min}$ and $m_{\chi}$, constraints can be placed on the fraction of DM taht PBHs compose as a function of the average mass of the binary black hole system. These constraints are shown in Fig. 27, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 are dark. Here, we limit ourselves to case (1) and case (3). These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm:2003vv, which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$. The parameters b=[1,2],r=[2,100]b=[-1,2],r=[2,100] determine the mass distribution of dark BHs. Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions. More observations of dark BHs would be needed to further narrow this region. However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
Figure 27: Taken from Singh:2020wiq. Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA. These plots assume an atomic model of DM that forms dark BHs. The constraints are obtained in three cases: (1) assuming none of the detected BHs are dark; (2) assuming only GW190425 is a dark BH; and (3) assuming both GW190425 and GW190814 are dark BHs. The latter two cases essentially arise from a joint posterior distribution conditioned on the rate densities inferred from the search results, which also constrain the minimum allowed mass in the assumed population distribution (see text for details). Even in the absence of any dark BHs, the constraints obtained in Singh:2020wiq significantly outperform existing microlensing experiments Macho:2000nvd; EROS-2:2006ryy; Niikura:2019kqi and those from dynamics of dwarf galaxies Brandt:2016aco; Koushiappas:2017chw. Additional constraints arise from PBH searches LIGOScientific:2019kan; Gaggero:2016dpq; Raidal:2017mfl; Raidal:2018bbj; Ali-Haimoud:2017rtz. SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.